Numerical Analysis I, Fall 2017 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 03

Due Monday, Oct 09. Goal:

1. Section 2.2: Problems 9, 14(a), 20.

2. Section 2.2: Problems 14(a). Find a suitable choice of α to accelerate the convergence using generalized fixed point iteration.

3. Section 2.3: Problems 13(a,b), 18, 17(c).